OpenStudy (anonymous):

A triangular pyramid has the same base and height as a triangular prism. The volume of the prism is 66 cubic inches. What is the volume of the triangular pyramid?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

the volume of the pyramid is 1/3 of the volume of the prism

OpenStudy (welshfella):

Volume of prism = area triangle * height Volume of pyramid = 1/3 * area triangle * height

OpenStudy (welshfella):

so volume of pyramid = 1/3 * volume of the prism = 1/3 * 66